STARK EFFECT STUDIES OF THE ELECTRONIC SPECTRUM OF 1-PHENYLPYRROLE AT HIGH RES-OLUTION. CHARGE TRANSFER STATES?^{a}

\underline{JESSICA A. THOMAS}, JUSTIN W. YOUNG, ADAM J. FLEISHER, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

We report Stark effect studies of the S_{1}-S_{0} electronic spectrum of 1-phenylpyrrole (1PP) at high resolution. Analyses of these data yield values of the magnitudes and orientations of the permanent electric dipole moments of both electronic states. When combined with an earlier analysis of the vibrationally resolved spectrum,^{b} the results provide new insights into the extent to which the S_{1} state of 1PP is a charge transfer state.

aWork supported by NSF CHE-0911117.

bJ.A. Thomas, et al., 64^{th} International Symposium on Molecular Spectroscopy, TB06, 2009.